# Testing tidal theory using Gaia binaries: the red giant branch

### v1

@@ 1 Introduction
In contrast, tidal theory has enjoyed some qualified success when applied to red giants. Verbunt & Phinney (1995) pioneered the study of tidal circularization in red giant stars. They showed that a set of $28$ binaries in open clusters are consistent with the theory of Zahn (1977), where the equilibrium tidal bulge is damped by eddy viscosity in a convective envelope: $6$ of the systems predicted by equilibrium tides to be circular are indeed so. This work was extended by Price-Whelan & Goodman (2018) to $234$ binaries with evolved primaries from the APOGEE survey. They also demonstrated that the orbital period within which all eccentricities are negligible (circularization period) increases with decreasing surface gravity (i.e., with advancing stellar evolution) in the primary star.
@@ 2.1 Sample Extraction
Now left with a total of $150,927$ binary systems, we further separate them by evolutionary status. As can be observed in the HR diagram of these systems (Fig. 1), the main-sequence binaries [^2] We will name a binary after the more luminous component, e.g., a MS binary, a giant binary. can be separated from the giant binaries by the following line:
$M_{G}={\rm Max}\{4.0,\,\left[(B_{p}-R_{p})_{\rm de-red}-1\right]^{7}\}\,$ (1)
with those brighter being giants and dimmer as main-sequences. This returns $119,425$ MS binaries and $31,502$ giant binaries.
@@ 2.2 Giant Binaries
Analogous to main-sequence stars, giant binaries also exhibit an upper envelope in the eccentricity-period space. As Fig. 4 shows, this upper envelope moves outwards in period as the primary star ascends the giant branch. The corresponding peri-centre period evolves from $\sim 6$ days on the main-sequence, to $\sim 200$ days in the red-clump stage. Such an evolution is clearly caused by events on the RGB, with tidal evolution being the most likely culprit.
@@
Figure 4: Eccentricity distributions of Gaia giant binaries, represented by kernel density estimates and separated into groups by the primary radii. In the left column the x-axis shows period, while in the right column we convert period to the ratio $a/R$. The eccentricity distribution is more sharply defined in the latter space. We measure the reach of tidal circularization in this space, following the same procedure as in Fig. 2. The shrinking arrows indicate both the eccentricity upper envelopes (teal curves, with constant values of $\beta=a(1-e)/R$, which largely follow the lowest contour at $0.05$) and the edges of the cool islands (vertical bars).
@@ 3.2 Governing equations
Here $\boldsymbol{\xi}$ is the Lagrangian displacement vector, which in a rigid rotator is related to the Eulerian velocity perturbation ${\bf v}$ by ${\bf v}=\partial_{t}\boldsymbol{\xi}.$ Meanwhile, $c_{s}^{2}=\text{d}P_{0}/\text{d}\rho_{0}$ gives the equilibrium sound speed, which deviates from the adiabatic sound speed ($c_{A}^{2}=\Gamma_{1}P_{0}/\rho_{0}$, where $\Gamma_{1}$ is the first adiabatic exponent) in any regions with stable (or unstable) entropy stratification. Additionally $\eta=\delta\rho/\rho_{0}$, $h=\delta P/\rho_{0}$, and $\delta\Phi$ describe Eulerian perturbations to the density, pressure, and gravitational potential, respectively. Ignoring non-adiabatic heat exchange, these variables satisfy the relations
@@ 3.3 Viscous dissipation
$\nu_{c}=\begin{cases}5\ell_{c}v_{c},&|\frac{\omega}{\omega_{c}}|<10^{-2}\\ \frac{1}{2}\ell_{c}v_{c}\left(\omega_{c}/\omega\right)^{1/2},&10^{-2}\leq|\frac{\omega}{\omega_{c}}|\leq 5\\ \frac{25}{\sqrt{20}}\ell_{c}v_{c}\left(\omega_{c}/\omega\right)^{2},&5\leq|\frac{\omega}{\omega_{c}}|\end{cases}$ (7)
@@
The tidal dissipation rate is given by the time-average of
$\mathcal{D}$ $=\int_{V^{*}}\rho_{0}\nu_{c}{\bf S}^{*}:{\bf S}\text{d}V,$ (8)
where the tensor contraction ${\bf S}^{*}:{\bf S}$ can be computed from the velocity field ${\bf v}=\partial_{t}\boldsymbol{\xi},$ with $\boldsymbol{\xi}$ in turn calculated as described in §3.4.
@@ 3.4 Methods for computing the tidal flow
Here $U_{mk}$ are coefficients that depend on the orbital semi-major axis $a$ and eccentricity $e$ (see Appendix C), $R$ is the (equatorial) radius of the tidally perturbed star, $Y_{2m}$ is a spherical harmonic of degree $2$ and order $m$, and both $m$ and $k$ are integers. The tidal frequency associated with each component of this Fourier decomposition is $\omega_{mk}=k\Omega_{0}-m\Omega$, where $\Omega_{0}=[G(M+M^{\prime})/a^{3}]^{1/2}$ is the orbital mean motion. This is the only place where we incorporate the effects of stellar rotation into our tidal calculations. For $e=0$, the only nonzero coefficients $U_{mk}$ are those with $m=k$. Eccentric orbits, on the other hand, involve driving by potentials with many more tidal frequencies.
We adopt a linear approximation, assuming that the overall tide can be described as the superposition of the (infinitely many) responses $\boldsymbol{\xi}_{mk}$ to individual potentials, $\boldsymbol{\xi}=\sum_{m,k}\boldsymbol{\xi}_{mk}$. Additionally we assume that each linear response takes the same time dependence as its driving tidal potential (so that $\partial_{t}\boldsymbol{\xi}_{mk}=-\text{i}\omega_{mk}\boldsymbol{\xi}_{mk}$ for each $m,k$). Each response and its time-averaged dissipation rate can then be computed independently. We omit the indices $m$ and $k$ unless needed.
@@ 3.5 Secular orbital evolution
[…]
$\hskip 10.00002pt3\left[\frac{2}{(1-e^{2})^{1/2}}-k\right]|X_{2k}|^{2}\kappa_{2k}-k|X_{0k}|^{2}\kappa_{0k}\Bigg{\}},$
@@
$\kappa_{mk}=\frac{8\pi G}{5R}\frac{\mathcal{D}_{mk}}{|U_{mk}|^{2}\omega_{mk}},$ (12)
where $\mathcal{D}_{mk}$ are the time-averaged dissipation rates associated with each linear response $\boldsymbol{\xi}_{mk}$ (see Equation 8). Note that $\text{sgn}[\kappa_{mk}]=\text{sgn}[\omega_{mk}]$ for positive definite dissipation (Ogilvie, 2013).
@@
$I\equiv\frac{1}{MR^{2}}\int_{V^{*}}\rho_{0}r^{2}\sin^{2}\theta\text{d}V\,.$ (14)
@@ 4.1 Single star evolution
Figure 5: Evolution of mass (top), radius (middle), and relevant frequencies (bottom) for a $1.5M_{\odot}$ star. The background colors indicate different evolutionary stages. Rapid radial expansions, accompanied by mass-loss, occur near the end of RGB and AGB (grey lines in the middle panel indicate the radial bins applied to our sample). Horizontal yellow lines indicate the Roche radius (middle panel) and tidal driving frequencies (lower panel, opacities scaled by amplitude for different $k$ values) for a $500$d, $e=0.5$ binary orbit (with a secondary mass of $1.0M_{\odot}$).
Another important distinction of giant stars is their turbulent state. A larger luminosity and lower mean density bring about faster convective velocities $v_{c}$, as the star expands and increases in scale height ($l_{c}$). In combination, we find that the convective frequency (with a mass-weighted average plotted in the lower panel of Fig. 6) remains largely constant throughout the star’s evolution, and is comparable to the orbital mean motion of a $\sim 500$ day orbit. This means for most orbits of concern ($100-1000$days), turbulent viscosity does not suffer strong suppression due to ‘fast tides.’ At the same time, we observe a dramatic rise in eddy viscosity during the RGB (and AGB), as is shown in Fig. 6. This rise is even more significant in the late stages, and leads to strong dissipation during those short intervals.
@@
Figure 7: Tidal quality factors (inversely related to the dissipative coefficients appearing in Equations 10-13), computed as functions of tidal period ($P_{t}=2\pi/\omega$), at the end of each evolutionary stage. Dashed lines indicate ‘equilibrium’ tidal calculations, while solid lines show ‘f-mode’ tidal calculations. The yellow lines indicate relevant tidal periods, as in the bottom panel of Fig. 5. The f-mode and equilibrium calculations disagree when the tidal frequency becomes comparable to the stellar dynamical frequency.
@@ 4.3 Initial binary orbits
Figure 8: Orbital evolution driven by f-mode tides. From top to bottom, the panels show orbital period, stellar rotation rate (relative to both the dynamical frequency and the orbital mean motion), and eccentricity for some representative initial conditions (differentiated by line colors). Orange circles indicate RLOF. Eccentricity damping takes place almost exclusively during radial expansion on the RGB and AGB. The spin evolution of a single star (yellow dashed line) matches spin evolution in the widest binaries (light-green color).
Figure 9: Evolution of orbital eccentricity (left panel) and primary spin (right panel, normalized by $\omega_{\rm dyn}$), as functions of orbital period, for an assortment of initial conditions (green tri-stars) and integrated with f-mode tides. Line colors indicate evolutionary stages, and orange outlined points indicate end points concluding in RLOF. The orange dashed line delineates the boundary of RLOF, in the absence of tides. After the RGB and the AGB, tidal circularization extends to binaries of periods $\sim 1000$ days and $\sim 2000$ days (respectively). In close binaries the primary star’s rotation synchronizes with the orbital mean motion (approaching $\Omega\sim 0.35\omega_{\rm dyn}$ just prior to RLOF). Tides in binaries with $P_{\text{orb}}\simeq 5000$ and $10^{4}$ days are likely to be strongly affected by spin-up to large fractions of $\Omega/\omega_{\rm dyn}$ during the RGB and AGB (respectively). At larger separations the spin evolution is dominated by mass loss.
@@ 4.4 Orbital evolution
We incorporate dissipation coefficients like those shown in Fig. 7 into the numerical integration of Equations (10)-(13), assuming that our $1.5M_{\odot}$ primary star is in an orbit with a $1M_{\odot}$ secondary. We draw initial eccentricities and periods from the Rayleigh and log-normal distributions described above. We additionally adopt an initial rotation rate of $\Omega=0.01\omega_{\rm dyn}$ in all cases, noting that this condition is quickly erased by spin evolution due to tides and mass loss.
We start each integration from the beginning of the RGB, and continue until the end of the AGB, or until the stellar radius surpasses the Roche radius ($(1-e)R_{\text{roche}}$ (with $R_{\text{roche}}$ given by Equation 15).
@@
Figure 10: Comparison of tidal calculations and Gaia giant binaries. The eccentricity-distance distributions of giant binaries are shown in the two left panel (kernal density estimates; left column for theory, middle for data with blue lines indicating mean errors), split by the primary radii, our proxy for age. The right column presents separation distributions. The bottom row singles out red clump stars, although they are also present in other rows as the second ‘cold island’ at larger separations. Overall, the theoretical calculations capture qualitative features of the observed eccentricity-distance distributions, but with quantitative disagreement at early stages of RGB.
@@ 4.5 Population-level Results
As the primary stars age and expand in radius, binaries with longer and longer orbital periods can be circularized by tides. (see Fig. 12). But the distributions of theoretically predicted binaries remain nearly fixed in $e$ vs. $a/R$ space. Fig. 10 shows that, regardless of primary radii, the upper envelope of eccentricity almost always remains at $\beta\equiv a(1-e)/R\approx 4$. At the same time, a cool island (of circular orbits) extends slightly beyond this, to $a/R\sim 6$. Such a simple description attests to the power of using $a/R$ as the ordinate. We also observe a second cold island at much larger distances. This second island is due to contamination by giants that have descended from the tip of their RGB (red-clump stars).
@@ 5 Comparison and Discussions
The Gaia giant binaries are already showcased in Fig. 4. Here we directly compare theory and data (Fig. 10), using f-mode tides as the default model. Fig. 11 summarizes our results in terms of critical $a/R$ ratios. While f-mode tides consistently produce eccentricity upper envelopes with $\beta\equiv a(1-e)/R\simeq 4$, the envelopes in the Gaia data evolve from $\beta\approx 6$ in the early RGB to $\beta\approx 3$ near the RGB tip. In addition, the data show cold islands of circular orbits extending to $a/R$ beyond $10$ in the early RGB phase, and to $\sim 6$ in the late phases. In contrast, f-mode tides produce cool islands that always extend to $a/R\approx 6$. This comparison leads us to conclude that an additional process is required to explain circularization in the early RGB, while f-mode tides largely suffice at late stages.
@@
One possibility is enhanced viscosity. It is possible that we have adopted a prescription for turbulent viscosity that is too conservative. During the RGB, the viscous prescription of Duguid et al. (2020) does cause some mild suppression due to ‘fast tides,’ in particular in zones deep below the stellar surface. And this suppression is stronger in early RGB, as the relevant binary orbits during this period are tighter (Fig. 4). It is unclear if removing this suppression would help reconcile the early RGB discrepancy.
The second possibility is internal gravity waves (Zahn, 1977). As Fig. 6 shows, the inner radiative cores in early RGB stars are more substantial in size. These cores are also highly stratified. Gravity waves that are tidally excited and propagate inward can break in these cores (Goodman & Dickson, 1998; Weinberg et al., 2017; Guo et al., 2023). This introduces another source of dissipation that is unaccounted for in our work.
@@
Two other ingredients, not considered here, may also affect the tides of early RGB binaries. One is related to rotation. Rotation introduces a new class of waves called ‘inertial waves’ in stars, and modifies internal gravity modes. The dissipation of inertial waves (Ogilvie & Lin, 2007; Ogilvie, 2014; Barker, 2022) and/or gravito-inertial modes (Xu & Lai, 2017) has been argued to play an important role in some cases. Inertial waves may enhance tidal dissipation in the early RGB, but not in the late RGB, because while both have extended convective envelopes, convective turbulence in the early stage is weaker and may not obliterate inertial-wave resonances. The development of differential rotation (here we assume the primary star rotates rigidly) may also affect tidal eccentricity damping in the early RGB.
@@
Lastly, there may be extra-tidal effects. For instance, RGB winds may be captured by a binary companion to form a circum-binary disk. This disk may modify the binary eccentricity. There are two arguments against this proposal. First, Rafikov (2016) pointed out that such disks are likely too low in mass to matter; second, such a disk, if it exists, is likely more prominent in late RGB phases (when winds are stronger) than the early phases. This may not help resolve the discrepancy we find here.
Figure 11: Extents of tidal circularization, obtained from Gaia binaries (teal color, from Fig. 4), and from our f-mode calculations (orange color, from Fig. 10), expressed in units of $a/R$ and plotted as functions of stellar radii (a proxy for evolutionary stage). Each extent (a shrinking arrow) is as defined in Fig. 2. While late RGB stars are well described by theory, early RGB stars appear to experience stronger dissipation than predicted by tidal theory. These latter stars also exhibit more extended cool islands than expected, reminiscent of the main-sequence phase.

### v2

@@ 1 Introduction
In contrast, tidal theory has enjoyed some qualified success when applied to red giants. Verbunt & Phinney (1995) pioneered the study of tidal circularization in red giant stars. They showed that a set of $28$ binaries in open clusters are consistent with the theory of Zahn (1977), where the equilibrium tidal bulge is damped by eddy viscosity in a convective envelope: $6$ of the systems predicted by equilibrium tides to be circular are indeed so. This work was extended by Price-Whelan & Goodman (2018) to $234$ binaries with evolved primaries from the APOGEE survey. They demonstrated that the orbital period within which all eccentricities are negligible (circularization period) increases with decreasing surface gravity (i.e., with advancing stellar evolution) in the primary star. Beck et al. (2018) and Beck et al. (2024) drew similar conclusions using giant binaries from the Kepler mission, and with seismically inferred ages (respectively).
@@ 2.1 Sample Extraction
Now left with a total of $150,927$ binary systems, we further separate them by evolutionary status. As can be observed in the HR diagram of these systems (Fig. 1), the main-sequence binaries [^2] We will name a binary after the more luminous component, e.g., an MS binary, a giant binary. can be separated from the giant binaries by the following line:
$M_{G}={\rm Min}[4.0,7(B_{p}-R_{p})_{\rm de-red}-3.5],$ (1)
with those brighter being giants and those dimmer being MSs. This returns $119,425$ MS binaries and $31,502$ giant binaries.
@@ 2.2 Giant Binaries
Analogous to main-sequence stars, giant binaries also exhibit an upper envelope in the eccentricity-period space. As Fig. 4 shows, this upper envelope moves outwards in period as the primary star ascends the giant branch. The corresponding peri-centre period evolves from $\sim 6$ days on the main-sequence, to $\sim 200$ days in the red-clump stage. Such an evolution (also apparent in the oscillating subsample identified by Beck et al., 2024) is clearly caused by events on the RGB, with tidal evolution being the most likely culprit.
@@
Figure 4: Eccentricity distributions of Gaia giant binaries, represented by kernel density estimates and separated into groups by the primary radii. In the left column the x-axis shows period, while in the right column we convert period to the ratio $a/R$. The eccentricity distribution is more sharply defined in the latter space. We measure the reach of tidal circularization in this space, following the same procedure as in Fig. 2. The shrinking arrows indicate both the eccentricity upper envelopes (teal curves, with constant values of $\beta=a(1-e)/R$ indicated by text annotations, which largely follow the lowest contour at $0.05$) and the edges of the cool islands (vertical bars).
@@ 3.2 Governing equations
Here $\boldsymbol{\xi}$ is the Lagrangian displacement vector, which in a rigid rotator is related to the Eulerian velocity perturbation ${\bf v}$ by ${\bf v}=\partial\boldsymbol{\xi}/\partial t.$ Meanwhile, $c_{s}^{2}=\text{d}P_{0}/\text{d}\rho_{0}$ gives the equilibrium sound speed, which deviates from the adiabatic sound speed ($c_{A}^{2}=\Gamma_{1}P_{0}/\rho_{0}$, where $\Gamma_{1}$ is the first adiabatic exponent) in any regions with stable (or unstable) entropy stratification. Additionally $\eta=\delta\rho/\rho_{0}$, $h=\delta P/\rho_{0}$, and $\delta\Phi$ describe Eulerian perturbations to the density, pressure, and gravitational potential, respectively. Ignoring non-adiabatic heat exchange, these variables satisfy the relations
@@ 3.3 Viscous dissipation
$\nu_{c}=\begin{cases}5\ell_{c}v_{c},&|\omega|/\omega_{c}<10^{-2}\\ \frac{1}{2}\ell_{c}v_{c}\left(\omega_{c}/|\omega|\right)^{1/2},&10^{-2}\leq|\omega|/\omega_{c}\leq 5\\ \frac{25}{\sqrt{20}}\ell_{c}v_{c}\left(\omega_{c}/|\omega|\right)^{2},&5<|\omega|/\omega_{c}\end{cases}$ (7)
@@
The volume-integrated rate of energy dissipation due to the effective viscosity $\nu_{c}$ is then
$\mathcal{D}$ $=\int_{V^{*}}2\rho_{0}\nu_{c}{\bf S}:{\bf S}\text{d}V.$ (8)
The tensor contraction ${\bf S}:{\bf S}$ can be computed from the velocity field ${\bf v}=\partial\boldsymbol{\xi}/\partial t,$ with $\boldsymbol{\xi}$ in turn calculated as described in §3.4.
@@ 3.4 Methods for computing the tidal flow
Here $U_{mk}$ are coefficients that depend on the orbital semi-major axis $a$ and eccentricity $e$ (see Appendix C), $R$ is the (equatorial) radius of the tidally perturbed star, $Y_{2m}$ is a spherical harmonic of degree $2$ and order $m$, and both $m$ and $k$ are integers. The tidal frequency associated with each component of this Fourier decomposition is $\omega_{mk}=k\Omega_{0}-m\Omega$, where $\Omega_{0}=[G(M+M^{\prime})/a^{3}]^{1/2}$ is the orbital mean motion. This frequency is the only place where we incorporate the effects of stellar rotation into our tidal calculations. For $e=0$, the only nonzero coefficients $U_{mk}$ are those with $m=k$. Eccentric orbits, on the other hand, involve driving by potentials with many more tidal frequencies.
We adopt a linear approximation, assuming that the overall tide can be described as the superposition of the (infinitely many) responses $\boldsymbol{\xi}_{mk}$ to individual potentials, $\boldsymbol{\xi}=\sum_{m,k}\boldsymbol{\xi}_{mk}$. Additionally we assume that each linear response takes the same time dependence as its driving tidal potential (so that $\partial\boldsymbol{\xi}_{mk}/\partial t=-\text{i}\omega_{mk}\boldsymbol{\xi}_{mk}$ for each $m,k$). Each response and its time-averaged dissipation rate can then be computed independently. We omit the indices $m$ and $k$ unless needed.
@@ 3.5 Secular orbital evolution
$\frac{\dot{e}}{e}$ $=\frac{(1-e^{2})}{e^{2}}\frac{\Omega_{o}\epsilon_{t}}{4}\sum_{k=-\infty}^{\infty}\Bigg{[}$ (11)
$\hskip 10.00002pt3\left(\frac{2}{(1-e^{2})^{1/2}}-k\right)|X_{2k}|^{2}\kappa_{2k}-k|X_{0k}|^{2}\kappa_{0k}\Bigg{]},$
@@
$\kappa_{mk}=\frac{8\pi G}{5R}\frac{\mathcal{D}_{mk}}{\omega_{mk}},$ (12)
where $\mathcal{D}_{mk}$ are the (amplitude-normalized and time-averaged) dissipation rates associated with each linear response $\boldsymbol{\xi}_{mk}$. Note that $\text{sgn}[\kappa_{mk}]=\text{sgn}[\omega_{mk}]$ for positive definite dissipation (Ogilvie, 2013).
@@
$I=\frac{1}{MR^{2}}\int_{V^{*}}\rho_{0}r^{2}\sin^{2}\theta\text{d}V\,.$ (14)
@@ 4.1 Single star evolution
Figure 5: Evolution of mass (top), radius (middle), and relevant frequencies (bottom) for a $1.5M_{\odot}$ star. The background colors indicate different evolutionary stages. Rapid radial expansions, accompanied by mass-loss, occur near the end of RGB and AGB (grey lines in the middle panel indicate the radial bins applied to our sample). Horizontal yellow lines indicate the Roche radius (middle panel) and tidal driving frequencies (lower panel, opacities scaled by tidal driving amplitudes) for a $500$d, $e=0.5$ binary orbit (with a secondary mass of $1.0M_{\odot}$).
Another important distinction of giant stars is their turbulent state. A larger luminosity and lower mean density bring about faster convective velocities $v_{c}$, as the star expands and increases in scale height ($l_{c}$). In combination, we find that the convective frequency (with a mass-weighted average plotted in the lower panel of Fig. 6) remains largely constant throughout the star’s evolution, and is comparable to the orbital mean motion of a $\sim 500$ day orbit. This means that for most orbits of concern ($100-1000$days), turbulent viscosity does not suffer strong suppression due to ‘fast tides.’ At the same time, we observe a dramatic rise in eddy viscosity during the RGB (and AGB), as is shown in Fig. 6. This rise is even more significant in the late stages, and leads to strong dissipation during those short intervals.
@@
Figure 7: Tidal quality factors (inversely related to the dissipative coefficients appearing in Equations 10-13), computed as functions of tidal period ($P_{t}=2\pi/\omega$), at the end of each evolutionary stage. Dashed lines indicate ‘equilibrium’ tidal calculations, while solid lines show ‘f-mode’ tidal calculations. The yellow lines indicate relevant tidal periods (and corresponding driving amplitudes), as in the bottom panel of Fig. 5. The f-mode and equilibrium calculations disagree when the tidal frequency becomes comparable to the stellar dynamical frequency.
@@ 4.3 Initial binary orbits
Figure 8: Orbital evolution driven by f-mode tides. From top to bottom, the panels show orbital period, stellar rotation rate (relative to both the dynamical frequency and the orbital mean motion), and eccentricity for some representative initial conditions (light/dark line colors indicate long/short orbital periods). Orange circles indicate RLOF. Eccentricity damping takes place almost exclusively during radial expansion on the RGB and AGB. The spin evolution of a single star (yellow dashed line) matches spin evolution in the widest binaries (light-green color).
Figure 9: Evolution of orbital eccentricity (left panel) and primary spin (right panel, normalized by $\omega_{\rm dyn}$), as functions of orbital period, for an assortment of initial conditions (green tri-stars) and integrated with f-mode tides. Line colors indicate evolutionary stages, and orange outlined points indicate end points concluding in RLOF. The orange dashed line delineates the boundary of RLOF, in the absence of tides. After the RGB and the AGB, tidal circularization extends to binaries of periods $\sim 1000$ days and $\sim 2000$ days (respectively). In close binaries the primary star’s rotation synchronizes with the orbital mean motion (approaching $\Omega\sim 0.35\omega_{\rm dyn}$ just prior to RLOF). Tides in binaries with $P_{\text{orb}}$ between $5000$ and $10^{4}$ days are likely to be strongly affected by spin-up to large fractions of $\Omega/\omega_{\rm dyn}$ during the RGB and AGB (respectively). At larger separations the spin evolution is dominated by mass loss.
@@ 4.4 Orbital evolution
We incorporate on-the-fly calculations of dissipation coefficients like those shown in Fig. 7 into the numerical integration of Equations (10)-(13), interpolating from oscillation mode properties pre-computed for each MESA snapshot (see Appendix B). We assume that our $1.5M_{\odot}$ primary star is in an orbit with a $1M_{\odot}$ secondary. We draw initial eccentricities and periods from the Rayleigh and log-normal distributions described above. We additionally adopt an initial rotation rate of $\Omega=0.01\omega_{\rm dyn}$ in all cases, noting that this condition is quickly erased by spin evolution due to tides and mass loss.
We start each integration from the beginning of the RGB, and continue until the end of the AGB, or until the stellar radius surpasses the Roche radius $(1-e)R_{\text{roche}}$ (with $R_{\text{roche}}$ given by Equation 15).
@@
Figure 10: Kernel density estimates computed from eccentricity and distance data for Gaia giant binaries (left) and theoretical tidal calculations (right). The panels are split vertically by the radius of the primary star (our proxy for age). The bottom row singles out red clump stars, although they are also present in other rows as the second “cool island” at larger separations. Overall, the theoretical calculations capture qualitative features of the observed eccentricity-distance distributions, but with quantitative disagreement during early stages of the RGB.
@@ 4.5 Population-level Results
As the primary stars age and expand in radius, binaries with longer and longer orbital periods can be circularized by tides. (see Fig. 12). But the distributions of theoretically predicted binaries remain nearly fixed in $e$ vs. $a/R$ space. Fig. 10 shows that, regardless of primary radii, the upper envelope of eccentricity almost always remains at $\beta\equiv a(1-e)/R\approx 4$. At the same time, a cool island (of circular orbits) extends slightly beyond this, to $a/R\sim 6$. Such a simple description attests to the power of using $a/R$ as the ordinate. We also observe a second cool island at much larger distances. This second island is due to contamination by giants that have descended from the tip of their RGB (red-clump stars).
@@ 5 Comparison and Discussions
The Gaia giant binaries are already showcased in Fig. 4. Here we directly compare theory and data (Fig. 10), using f-mode tides as the default model. Fig. 11 summarizes our results in terms of critical $a/R$ ratios. While f-mode tides consistently produce eccentricity upper envelopes with $\beta\equiv a(1-e)/R\simeq 4$, the envelopes in the Gaia data evolve from $\beta\approx 6$ in the early RGB to $\beta\approx 3$ near the RGB tip. In addition, the data show cool islands of circular orbits extending to $a/R$ beyond $10$ in the early RGB phase, and to $\sim 6$ in the late phases. In contrast, f-mode tides produce cool islands that always extend to $a/R\approx 6$. This comparison leads us to conclude that an additional process is required to explain circularization in the early RGB, while f-mode tides largely suffice at late stages.
@@
One possibility is enhanced viscosity. It is possible that we have adopted a prescription for turbulent viscosity that is too conservative. During the RGB, the viscous prescription of Duguid et al. (2020) does cause some mild suppression due to ‘fast tides,’ in particular in zones deep below the stellar surface. And this suppression is stronger in the early RGB, as the relevant binary orbits during this period are tighter (Fig. 4). It is unclear if a different viscous prescription would help reconcile the early RGB discrepancy.
The second possibility is internal gravity waves (Zahn, 1977). As Fig. 6 shows, the inner radiative cores in early RGB stars are more substantial in size. These cores are also highly stratified. Gravity waves that are tidally excited and propagate inward can break in these cores (Goodman & Dickson, 1998; Weinberg et al., 2017; Guo et al., 2023; Esseldeurs et al., 2024). This introduces another source of dissipation that is unaccounted for in our work.
@@
Two other ingredients, not considered here, may also affect the tides of early RGB binaries. One is related to rotation. Rotation introduces a new class of waves called ‘inertial waves’ in stars, and modifies internal gravity modes. The dissipation of inertial waves (Ogilvie & Lin, 2007; Ogilvie, 2014; Barker, 2022) and/or gravito-inertial modes (Xu & Lai, 2017) has been argued to play an important role in some cases. Inertial waves may enhance tidal dissipation in the early RGB, but not in the late RGB, because while both have extended convective envelopes, convective turbulence in the early stage is weaker and may not obliterate inertial-wave resonances (but see Beck et al., 2018, who predict that inertial waves should have a limited impact on the RGB due to a reduced radiative core size). The development of differential rotation (here we assume the primary star rotates rigidly) may also affect tidal eccentricity damping in the early RGB.
@@
Lastly, there may be extra-tidal effects. For instance, RGB winds may be captured by a binary companion to form a circum-binary disk. This disk may modify the binary eccentricity. There are two arguments against this proposal. First, Rafikov (2016) pointed out that such disks are likely too low in mass to matter; second, such a disk, if it exists, is likely more prominent in late RGB phases (when winds are stronger) than early phases. This may not help resolve the discrepancy we find here.
Figure 11: Extents of tidal circularization, obtained from Gaia binaries (teal color, from Fig. 4), and from our f-mode calculations (orange color, from Fig. 10), expressed in units of $a/R$ and plotted as functions of stellar radii (a proxy for evolutionary stage). The extent of each shrinking arrow is defined in Fig. 2. While late RGB stars are well described by tidal theory, early RGB stars appear to experience stronger dissipation than predicted. These latter stars also exhibit more extended cool islands than expected, reminiscent of the main-sequence phase.